0.10 mol of argon gas is admitted to an evacuated 60 cm3 container at 20∘C. The gas then undergoes an isobaric heating to a temperature of 350 ∘C.

What is the final volume of the gas?
Answer #1

Apply ideal gas law
p·V = N·R·T

The initial pressure after filling is
p = N·R·T/V

Next u raise temperature to T' . Volume changes to V' while N and p remain unchanged.
V' = N·R·T' / p
replace p by the first expression
V' = N·R·T' / (N·R·T/V)
     = V·T'/T
     = 60cm³ · 623K / 293K
     = 127.6 cm³